OpenStudy (anonymous):

What is the product of the radical expression? (7-√2)(8+√2) A. 54 + 56 √2 B. 13 + 15 √2 C. 54 - √2 D. 58 + 56 √2 Please help :)

OpenStudy (jadedry):

What does \[(x-y)(z+y) = ?\] If you simplify that you will be able to get the answer to this question.

OpenStudy (anonymous):

first we multiply: \[(7-\sqrt{2}) \times (8 + \sqrt{2})\] then we get: \[56 - \sqrt{2} -2\] which simplifies to \[54 - \sqrt{2}\] which happens to be C